Yonnie:

$A is divided in the ratio of x:y. What is the difference, in dollars, between the two amounts if x>y?

sillybilly123:

Divide A in proportions: \(\alpha = A \cdot \dfrac{x}{x+y} \) AND \( \beta = A \cdot \dfrac{y}{x+y} \) To first double check: \(\alpha + \beta = A ~ \checkmark\) and \(\dfrac{\alpha}{\beta} = \dfrac{x}{y} ~ \checkmark\) then work out: \(\alpha - \beta\)